Barium ion with a 2+ charge travels at a speed 800 m/s and is acted upon by a magnetic force having a magnitude of 1700.0 N. Determine the magnitude of magnetic field acting perpendicularly on Ba^(2+).


1
Expert's answer
2022-05-20T13:19:46-0400
"F=qvB\sin\theta"

"B=\frac{F}{qv\sin\theta}\\
=\frac{1700}{2*1.6*10^{-19}*800*\sin90^\circ}=6.64*10^{18}\:\rm T"